How many litres of a 30% acid solution should be added to 20 litres of a 60% acid solution in order to obtain a 50% acid solution?
Answer by mananth(16946)

------- percent ---------------- quantity
Alcohol I 60 ---------------- 20 ounces
Alcohol II 30 ---------------- x ounces
Total 50 ---------------- 20 + x ounces

60*20+30 *x=50(20 +x )

1200+30 x=1000+50x

30 x - 50 x = 1000 - 1200

-20 x = -200 / -20
x = 10 ounces 30 % Alcohol II